# When Reasoning Meets Compression: Benchmarking Compressed Large Reasoning Models on Complex Reasoning Tasks

### v1

@@ Abstract
Recent open-source large reasoning models (LRMs) exhibit strong performance on complex reasoning tasks, but their large parameter count makes them prohibitively expensive for individuals. The compression of large language models (LLMs) offers an effective solution to reduce cost of computational resources. However, systematic studies on the performance of compressed LLMs in complex reasoning tasks, especially for LRMs, are lacking. Most works on quantization and pruning focus on preserving language modeling performance, while existing distillation works do not comprehensively benchmark student models based on reasoning difficulty or compression impact on knowledge and reasoning. In this paper, we benchmark compressed DeepSeek-R1 models on four different reasoning datasets (AIME 2024, FOLIO, Temporal Sequences of BIG-Bench Hard, and MuSiQue), ranging from mathematical to multihop reasoning, using quantization, distillation, and pruning methods. We benchmark 2.51-, 1.73-, and 1.58-bit R1 models that adopt dynamic quantization. We also benchmark distilled R1 models that are based on LLaMA or Qwen and run SparseGPT on them to obtain various sparsity levels. Studying the performance and behavior of compressed LRMs, we report their performance scores and test-time compute (number of tokens spent on each question). Notably, using MuSiQue, we find that parameter count has a much greater impact on LRMs’ knowledge memorization than on their reasoning capability, which can inform the choice of compression techniques. Through our empirical analysis of test-time compute, we find that shorter model outputs generally achieve better performance than longer ones across several benchmarks for both R1 and its compressed variants, highlighting the need for more concise reasoning chains.
@@ 1 Introduction
Large reasoning models (LRMs) excel at complex reasoning tasks. DeepSeek-R1 (Guo et al., 2025) is the first open-source LRM to successfully replicate OpenAI’s O1-level performance. Trained based on DeepSeek-V3 (DeepSeek-AI et al., 2025), R1 adopts large-scale reinforcement learning without supervised fine-tuning (SFT). Its open-source nature facilitates research on reasoning models (Srivastava et al., 2025; Mondillo et al., 2025). However, due to its large size (671B total parameters), deploying it locally can be costly and even infeasible for individuals, which hinders AI democratization.
Compression of large language models (LLMs) includes quantization, distillation, and pruning. Its goal is to reduce computational resources (e.g., GPU memory and disk space) and provide inference speedup. As the weights of vanilla LLMs are typically in 16-bit floating values (e.g.., FP16 or BF16), quantization means to convert these high-precision values to lower-precision ones. Representative techniques include dynamic quantization by Unsloth (Daniel Han & team, 2023), activation-aware quantization (Lin et al., 2024), and post-training quantization (Frantar et al., 2022). Distillation aims to transfer knowledge from a larger teacher model to a smaller student model under black-box (Li et al., 2024a) or white-box (Gu et al., 2024) settings. As for pruning, it aims to sparsify LLMs by introducing 0s to replace nonzero weights. Representative techniques include unstructured pruning (Zhang et al., 2024a; Frantar & Alistarh, 2023) and structured pruning (Xia et al., 2024; Ma et al., 2023).
Although existing quantization and pruning methods claim to preserve the performance of non-reasoning LLMs after compression such as LLaMA (Touvron et al., 2023) and OPT (Zhang et al., 2022)), exploring these methods on LRMs offers a timely research of democratizing the latest LRMs. Moreover, since these methods mainly adopt perplexity and simple end-task scores for evaluation, the assessment of compressed LRMs will be fundamentally different. Without focusing on basic linguistic capabilities (e.g., simple language understanding and generation), evaluation of compressed reasoning models needs to align with their uncompressed counterparts. Regarding distillation, recent works either fail to comprehensively evaluate their student models on diverse reasoning benchmarks of varying difficulty or neglect to consider compression impact on knowledge and reasoning. A comprehensive evaluation that includes all popular compression methods is valuable.
Therefore, due to the lack of compression research on LRMs, we benchmark the performance and behavior of compressed DeepSeek-R1 on various reasoning tasks in this paper. Our benchmarking framework is shown in Figure 1. We test dynamic quantization (Daniel Han & team, 2023), distillation with SFT (Guo et al., 2025), and SpraseGPT (Frantar & Alistarh, 2023) on R1 (or distilled R1) due to their popularity and effectiveness. Specifically, our chosen dynamic quantization reduces the MoE (Mixture of Experts) layers of R1 model to 2.51-bit, 1.73-bit, and 1.58-bit precision. Testing distillation with SFT, we evaluate the distilled R1 models that adopt LLaMA-70B (Meta AI, ) and Qwen-32B (Qwen et al., 2025) as students. Using SparseGPT, we prune the distilled R1 models to different sparsity levels. We select four complex reasoning benchmarks: mathematical, logical, temporal, and multihop reasoning.
Figure 1: An overview of selected compression methods and tasks. We evaluate R1 and its compressed variants based on their performance scores and test-time compute.
In addition to assessing model performance on selected benchmarks, we study the behavior of compressed R1 models by measuring their test-time compute (number of tokens spent on each question). Specifically, we compare their performance on the shortest and longest outputs. Our major findings are:
@@
  • Parameter count has a much greater impact on LRMs’ knowledge memorization than on their reasoning capability. Therefore, both distillation and pruning are discouraged when the end task relies on models’ parametric knowledge. Otherwise, larger student models or low sparsity levels are recommended.
  • In almost all cases, both R1 and its compressed variants achieve higher scores when they spend less compute during inference time. Shorter outputs tend to be more accurate than longer ones across our benchmarks. This highlights the need to reduce verbosity to improve reasoning performance.
2 Related Work
2.1 Quantization
PTQ is more popular in terms of the number of recent publications, because there is no retraining involved. For example, GPTQ (Frantar et al., 2022) is a one-shot weight quantization method that uses approximate second-order information, while AWQ (Lin et al., 2024) leverages activation distribution for finding the salient weight channels to skip. Other PTQ methods include weight-activation quantization (Shao et al., 2024; Yao et al., 2022; Liu et al., 2023) and KV cache quantization (Hooper et al., 2024; Liu et al., 2024b).
2.2 Distillation
Distillation involves two settings: black-box and white-box settings. For black-box setting, teacher model is typically a closed-source LLM and only the outputs of teacher are available for student model. For white-box setting, both weights and output distribution of the teacher model are available. Existing black-box distillation (Huang et al., 2024; Li et al., 2024b; Ho et al., 2023; Huang et al., 2022; Li et al., 2024a) prompts the teacher model to generate a training dataset for the student to learn. Specifically, researchers have started to distill OpenAI’s O1 model (Huang et al., 2024), which marks the beginning of LRMs compression. White-box distillation allows the student model to learn from teacher’s knowledge representation. Works has been done to align the output distribution (Agarwal et al., 2024; Gu et al., 2024) or the hidden representation (Liang et al., 2023) between teacher and student models.
2.3 Pruning
2.4 LRMs
3 Background
Scope We benchmark quantization, distillation, and pruning methodologies, because we consider them as major LLMs compression paradigms in this paper. Due to lower popularity, other methodologies that can potentially facilitate compression such as low-rank factorization (Sharma et al., 2024; Srebro & Jaakkola, 2003) are not elaborated. When selecting models for compression, we focus exclusively on LRMs.
Recent efforts on LRMs compression As reviewed in Section 2, compression on LRMs (not LLMs) is relatively underexplored. Recently, Unsloth (Daniel Han & team, 2023) introduces dynamic quantization by dynamically opting not to quantize certain LLM weights. By combining four ideas (Yu et al., 2024; Dettmers et al., 2023; Ma et al., 2024; Gerganov, 2023), Unsloth quantizes DeepSeek-R1 into four variants: 2.51-, 2.22-, 1.73-, and 1.58-bit models. Here the number of bits (e.g., 2.51-bit) represents the precision of MoE weights, which make up 88% of the total R1 weights. DeepSeek-R1 (Guo et al., 2025) also comes with several distilled models via black-box distillation.
@@ 3 Background
Current bottlenecks Very few quantization or pruning works have demonstrated effectiveness on LRMs, as LRMs have only recently emerged. Moreover, evaluation of quantized or pruned LRMs will be different from quantized or pruned LLMs. Current works evaluate quantization and pruning performance primarily using perplexity and simple end tasks, such as the EleutherAI evaluation harness (Gao et al., 2024) and commonsense reasoning. However, compressed LRMs should be assessed on more complex reasoning tasks with varying difficulty levels. As for distillation, although recent distillation works tend to test on more challenging reasoning tasks (compared to other compression literature) such as GSM8K (Cobbe et al., 2021), it is unclear how the compression of LRMs affects models’ parametric knowledge and reasoning capability. Some of them do not comprehensively select diverse reasoning benchmarks. Our benchmarking framework aims to address these bottlenecks.
#Questions #Pass Answer Type Metric Knowledge required?
AIME 2024 30 1 Integer Accuracy False
FOLIO 203 1 True/False/Uncertain Accuracy False
Temporal 250 1 (A)/(B)/(C)/(D) Accuracy False
MuSiQue 100 1 A few words (EM, F1) True
Table 1: Statistics of selected benchmarks: AIME 2024, FOLIO, Temporal Sequences of BIG-Bench Hard, and MuSiQue. We use exact match (EM) and F1 to measure model performance on MuSiQue.
4 Experiment Setup
4.1 Reasoning Benchmarks
To address the evaluation bottleneck discussed in Section 3, we select four different reasoning benchmarks with varying levels of difficulty: AIME 2024 (Mathematical Association of America, ) for mathematical reasoning, FOLIO (Han et al., 2024) for logical reasoning, Temporal Sequences of BIG-Bench Hard (Suzgun et al., 2022) for temporal reasoning, and MuSiQue (Trivedi et al., 2022) for multihop reasoning. Table 1 summarizes their statistics. For each model, we do a single pass on every benchmark.
Unlike other benchmarks where we select all questions, we randomly sample 100 out of 1000 from MuSiQue, as it is much larger. This pool of 1000 questions is the same as SiReRAG (Zhang et al., 2025) and HippoRAG (Jimenez Gutierrez et al., 2024). Since MuSiQue requires knowledge memorization besides multihop reasoning, it is popular under retrieval-augmented generation (RAG) setup. Instead of RAG, we follow a closed-book setting (directly prompting LRMs to get final answers) to test models’ parametric knowledge. In this case, unlike other benchmarks that test only reasoning capability, we use MuSiQue to evaluate both reasoning and knowledge memorization. Additional details of selected benchmarks are specified in Appendix A.
4.2 Selected Compression Methods
We decide to compress DeepSeek-R1 due to it large parameter count. For quantization, we select 2.51-, 1.73-, and 1.58-bit models [^1] https://huggingface.co/unsloth/DeepSeek-R1-GGUF by Unsloth. These dynamically quantized models are among the most widely used quantization methods on R1. As for distillation, we select the two largest distilled models that accompany R1: DeepSeek-R1-Distill-Llama-70B [^2] https://huggingface.co/deepseek-ai/DeepSeek-R1-Distill-Llama-70B and DeepSeek-R1-Distill-Qwen-32B [^3] https://huggingface.co/deepseek-ai/DeepSeek-R1-Distill-Qwen-32B. Since we do not find a pruning method designed for models with around 700B parameters and MoE layers (e.g., SparseGPT can only prune models with around 176B parameters), we use SparseGPT to prune our selected distilled models. In order to precisely analyze the collapse point of this pruning method, we run SparseGPT on each distilled model seven times to get 10%, 30%, 40%, 50%, 60%, 70%, and 80% sparsity levels.
4.3 Evaluation Metrics
Accuracy metric is used for AIME 2024, FOLIO, and Temporal Sequences. Since MuSiQue involves question answering and its answers are in a few words, we adopt exact match (EM) and F1 to measure question answering performance of different models. Without worrying inference speedup of selected compression methods, we focus on benchmarking their performance. The reason is that these methods run on different inference platforms. As a result, different optimization strategies might be adopted when we do inference, and it is hard to control the consistency of inference optimization across various platforms.
Our analysis of test-time compute involves computing the number of tokens spent on each question. We first preprocess each output by extracting reasoning tokens and sentences of final model prediction. Instead of relying on a specific tokenizer, we then treat words within each output as tokens for simplicity. We study whether there is any relationship between length and reasoning performance.
4.4 Implementation Details
Required by our quantized models, we run their inference on llama.cpp [^4] https://github.com/ggml-org/llama.cpp. While our pruned and distilled models can be deployed on various inference platforms, we use vLLM (Kwon et al., 2023) for its fast inference. In order to comprehensively analyze performance change after compression, we also evaluate R1 on our reasoning benchmarks by using DeepSeek API. Aligning with DeepSeek-R1 report (Guo et al., 2025), we keep the same parameters for all models: maximum generation length is set to $3276832768$, temperature is set to $0.60.6$, and top-p value is set to $0.950.95$.
5 Results and Analysis
Our results and analysis aim to answer the following research questions:
  • RQ 1: How does each compression methodology compare against each other (5.1)?
  • RQ 2: Do our selected compression methodologies collapse (5.2)?
  • RQ 3: What is the recommended methodology to compress LRMs (5.1, 5.3, and 5.5)?
  • RQ 4: Is it a good sign when models spend extra amounts of test-time compute on certain questions (5.4)?

### v2

# When Reasoning Meets Compression: Understanding the Effects of LLMs Compression on Large Reasoning Models

@@ Abstract
Compression methods, including quantization, distillation, and pruning, improve the computational efficiency of large reasoning models (LRMs). However, existing studies either fail to sufficiently compare all three compression methods on LRMs or lack in-depth interpretation analysis. In this paper, we investigate how the reasoning capabilities of LRMs are compromised during compression, through performance benchmarking and mechanistic interpretation. To uncover the effects of compression on reasoning performance, we benchmark quantized, distilled, and pruned DeepSeek-R1 models on four reasoning datasets (AIME 2024, FOLIO, Temporal Sequences, and MuSiQue). To precisely locate compression effects on model weights, we adapt difference of means and attribution patching techniques, focusing on the activation of every linear component in compressed LRMs, to interpret fine-grained causal relationships between weights and various reasoning capabilities. This fine-grained interpretation addresses a fundamental question of compression: which weights are the most important for reasoning? Overall, we find dynamically quantized 2.51-bit R1 reaches close-to-R1 performance. With empirical verification, we present three main findings that generalize across both Llama and Qwen: (1) Weight count has a greater impact on LRMs’ knowledge memorization than reasoning, highlighting the risks of pruning and distillation; (2) The MLP up projection in the final layer of distilled LRMs is one of the most important components, offering a new perspective on locating critical weights — a fundamental problem in model compression; and (3) Current quantization methods overly compress the final-layer modules and MLP gate projections, so protecting just 2% of all weights that are excessively compressed can raise average accuracy by 6.57%, greatly surpassing the state-of-the-art.
@@ 1 Introduction
Large reasoning models (LRMs) such as DeepSeek-R1 (Guo et al., 2025) excel at complex reasoning tasks. However, due to their large sizes, deploying them can be costly and even infeasible for individuals, which hinders AI democratization. Compression methods including quantization, distillation, and pruning reduce computational resources (e.g., GPU memory and disk space). Representative quantization techniques include dynamic quantization by Unsloth (Daniel Han & team, 2023), activation-aware quantization AWQ (Lin et al., 2024), and post-training quantization GPTQ (Frantar et al., 2022). Current distillation involves black-box (Li et al., 2024a) or white-box (Gu et al., 2024) settings. Representative pruning techniques include unstructured (Zhang et al., 2024a; Frantar & Alistarh, 2023) and structured pruning (Xia et al., 2024; Ma et al., 2023).
However, existing works do not sufficiently study the performance of compression method on LRMs (Liu et al., 2025; Srivastava et al., 2025; Feng et al., 2025). Although current quantization and pruning methods claim to preserve the performance of general-purpose LLMs, benchmarking both of them on LRMs with more reasoning-intensive datasets helps compare their collapse point. Regarding distillation, recent works either fail to comprehensively evaluate their student models on diverse reasoning benchmarks of varying difficulty or neglect to consider distillation effect on knowledge and reasoning (Huang et al., 2024; Agarwal et al., 2024). Another research gap is the lack of interpretability of compression effects on LRMs. It is necessary to interpret how compression methods affect LRMs, as such analysis can reveal existing bottlenecks and provide guidance for future compression research.
Figure 1: An overview of our pipeline. Left: We benchmark compressed R1 variants on various reasoning tasks. Right: By computing weight importance towards a specific reasoning behavior (a dot product of the steering vector and gradients with respect to an LRM’s activations), we study the compression effects on individual weight. We empirically verify our findings on weight importance by selectively quantizing or protecting a module to test its importance.
Therefore, due to the lack of compression works on LRMs, we study this fundamental research question: How are the reasoning capabilities of LRMs compressed during compression? We answer it from two perspectives — performance benchmarking and mechanistic interpretation. We first benchmark compressed DeepSeek-R1 on various reasoning tasks to investigate how model compression affects performance. We test dynamic quantization (Daniel Han & team, 2023), distillation with supervised fine-tuning (SFT) (Guo et al., 2025), SparseGPT (Frantar & Alistarh, 2023), AWQ (Lin et al., 2024), GPTQ (Frantar et al., 2022), GPTAQ (Li et al., 2025), and ANY4/3 (Elhoushi & Johnson, 2025) on R1 (or distilled R1). Then, we apply mechanistic interpretability to quantify weight contribution towards four core reasoning capabilities of LRMs: backtracking, uncertainty estimation, example testing, and adding knowledge. By focusing on the activation of every linear component in compressed LRMs, we adapt difference of means (Arditi et al., 2024) to extract steering vectors and attribution patching (Syed et al., 2023) to compute weight importance. Unlike previous analysis (Venhoff et al., 2025) that only measures layer-wise weight contribution, our weight importance scores offer more fine-grained interpretation of weight contribution, addressing the fundamental compression question of locating important weights. By comparing weight importance scores between compressed LRMs and original LRMs, we quantify the effects of distillation and quantization on model weights [^1] Our interpretation code is at https://github.com/psunlpgroup/Compression-Effects.. Our analysis framework is shown in Figure 1.
With empirical verification, our key findings are summarized below:
@@
  • Weight count has a greater impact on LRMs’ knowledge memorization than their reasoning capabilities, highlighting the compression effects of pruning and distillation. Thus, both distillation and pruning are discouraged when tasks require LRMs’ parametric knowledge.
  • The mlp.up_proj in the final layer of R1 distilled models emerges as one of the most important model components, addressing a core concern in pruning and quantization literature: identifying critical weights. Quantizing only this matrix to 3-bit reduces the average accuracy by 16.3%.
  • Final-layer modules, along with the mlp.gate_proj of R1 distilled Llama and Qwen, are overly compressed by popular quantization methods, highlighting the need for greater attention to preserving their weight precision. A successful protection of only final-layer MLP modules could raise average accuracy by 6.57%, with gains of up to 23.17% over the state-of-the-art quantization.
2 Problem Formulation
2.1 Background
As discussed in Section 1, compression on LRMs (not LLMs) is relatively underexplored. We conduct a thorough literature review in Appendix B.
@@ 2.1 Background
Bottlenecks on evaluation. Few quantization or pruning methodologies have sufficiently demonstrated effectiveness on LRMs. Current works evaluate quantization and pruning performance primarily using perplexity and simple end tasks, such as the EleutherAI evaluation harness (Gao et al., 2024) and commonsense reasoning. However, quantized or pruned LRMs should be assessed on more complex reasoning tasks with varying difficulty levels. For distillation, although recent works tend to test on more challenging reasoning tasks (compared to other compression literature) such as GSM8K (Cobbe et al., 2021), it is unclear how the compression of LRMs affects models’ parametric knowledge and reasoning capability. Some of them do not comprehensively select diverse reasoning benchmarks (Agarwal et al., 2024). Our benchmarking aims to address these bottlenecks.
Bottlenecks on in-depth analysis. The lack of interpretability of compression effects on LRMs is a key bottleneck of in-depth analysis for compressed LRMs. Being able to interpret the difference between original and compressed LRMs offers a new way to analyze the effects of compression. As a result, better compression approaches can be developed. A recent work (Venhoff et al., 2025) interprets several R1 distilled LRMs, but their focus is not on understanding compression effects.
Recent efforts. Recent benchmarking (Liu et al., 2025) and survey (Feng et al., 2025; Srivastava et al., 2025) papers have begun to evaluate compressed LRMs on more complex reasoning datasets, but they all lack in-depth interpretation of compression effects and do not comprehensively compare different compression strategies. As for compressed LRMs, Unsloth (Daniel Han & team, 2023) introduces dynamic quantization by dynamically opting not to quantize certain LLM weights. DeepSeek-R1 (Guo et al., 2025) also comes with several distilled models via black-box distillation. Our interpretation analysis aims to demystify the effects of LLMs compression on LRMs, providing a systematic understanding of existing compressed LRMs.
2.2 Mechanistic Interpretation
For our interpretation analysis, we target four core reasoning behaviors following an existing work (Venhoff et al., 2025): backtracking, uncertainty estimation, example testing, and adding knowledge. We prompt GPT-4o to locate token sequences of each behavior from the output tokens of our LRMs. To interpret different compression strategies, we adapt difference of means and attribution patching by computing the activation of every linear module in each layer. This allows us to compute the causal relationship between each weight matrix and our target reasoning behaviors.
Difference of Means. To compute the numerical representation in activation space of each reasoning behavior, we adapt difference of means method (Venhoff et al., 2025; Arditi et al., 2024) to extract the steering vector $\mathbf{u}_{m\ell}^{c}$ for each linear module $m$ at layer $\ell$ for behavior $c$:
$\mathbf{u}_{m\ell}^{c}=\frac{1}{\lvert\mathcal{D}_{+}\rvert}\sum_{s_{i}^{c}\in\mathcal{D}_{+}}\overline{\mathbf{a}}_{m\ell}^{c}(s_{i}^{c})-\frac{1}{\lvert\mathcal{D}_{-}\rvert}\sum_{s_{j}\in\mathcal{D}_{-}}\overline{\mathbf{a}}_{m\ell}(s_{j}),\quad\text{with}\quad\overline{\mathbf{a}}_{m\ell}^{c}(s_{i}^{c})=\frac{1}{\lvert s_{i}^{c}\rvert}\sum_{t\in s_{i}^{c}}\mathbf{a}_{m\ell}(t)$
where $s_{i}^{c}$ denotes the token sequence corresponding to a specific reasoning behavior $c$ along with its five preceding tokens as output by an LRM, $s_{j}$ is the token sequence of the entire LRM output (prompt and output tokens), $\mathcal{D}_{+}$ is the set of output instances containing at least one token sequence labeled with $c$, $\mathcal{D}_{-}$ is the set of all output instances, $\mathbf{a}_{m\ell}(t)$ is the activation of module $m$ at layer $\ell$ at token $t$, $\overline{\mathbf{a}}_{m\ell}^{c}(s_{i}^{c})$ is the average of $\mathbf{a}_{m\ell}(t)$ across all tokens in $s_{i}^{c}$, and similarly, $\overline{\mathbf{a}}_{m\ell}(s_{j})$ is the average of $\mathbf{a}_{m\ell}(t)$ across all tokens in $s_{j}$. We then normalize $\mathbf{u}_{m\ell}^{c}$ to $\tilde{\mathbf{u}}_{m\ell}^{c}$: $\tilde{\mathbf{u}}_{m\ell}^{c}=\mathbf{u}_{m\ell}^{c}\cdot\frac{\left\lVert\overline{\mathbf{a}}_{m\ell}^{\mathrm{all}}\right\rVert_{2}}{\left\lVert\mathbf{u}_{m\ell}^{c}\right\rVert_{2}}$ where $\overline{\mathbf{a}}_{m\ell}^{\mathrm{all}}$ denotes the mean activation across all tokens in $\mathcal{D}_{-}$.
Attribution Patching. To find the causally relevant LRMs components with respect to each reasoning behavior, we adapt attribution patching (Syed et al., 2023) method to compute the importance score $\mathbf{I}_{m\ell}^{c}$ of each linear module.
$\mathbf{I}_{m\ell}^{c}\approx\frac{1}{\lvert\mathcal{D}_{+}\rvert}\left|\sum_{s_{i}^{c}\in\mathcal{D}_{+}}\big(\tilde{\mathbf{u}}_{m\ell}^{\,c}\big)^{\top}\,\frac{\partial}{\partial\mathbf{a}_{m\ell}}\,\mathcal{L}\!\left(s_{i}^{c}\right)\right|$
where $\mathcal{L}\!\left(s_{i}^{c}\right)$ is the cross-entropy loss of $s_{i}^{c}$. A higher $\mathbf{I}_{m\ell}^{c}$ means a stronger causal relationship between $c$ and the linear module $m$ at layer $\ell$, helping us locate the most important weights responsible for reasoning capabilities (a fundamental problem for quantization and pruning works).
2.3 Decoding Compression Effects
To decode compression effects, we compute the relative importance $\mathbf{RI}_{m\ell}^{c}$ of each weight matrix ($\mathbf{I}_{m\ell}^{c}$ divided by $\sum_{m}\sum_{\ell}\mathbf{I}_{m\ell}^{c}$) and track how it changes because of compression (importance shift). Specifically, we measure the change of $\mathbf{RI}_{m\ell}^{c}$ from R1 distilled Llama-8B to original meta-llama/Llama-3.1-8B to understand distillation effect (Section 4). Likewise, the importance shift from the R1 distilled models to their quantized versions indicates quantization effect (Section 5). For the R1 distilled models, we also compute the $\mathbf{I}_{m\ell}^{c}$ of their weights to complement our findings on the distillation effect.
We hypothesize that the importance shift should be minimal in the ideal case, as a compressed LRM should remain as close as possible to its original counterpart (the more reasoning-capable model). When visualizing the importance shift from an LRM to its compressed variant (or from a distilled model to its backbone), we only consider decreases in $\mathbf{RI}_{m\ell}^{c}$. By definition, the relative importance of each weight matrix is normalized to sum to one, so any increase in relative importance necessarily compensates for decreases elsewhere. Since it is more informative to track cases where the $\mathbf{RI}_{m\ell}^{c}$ of a more reasoning-capable model decreases (e.g., when the reasoning capability of a weight matrix is diminished), we set all increases in relative importance to zero.
2.4 Scope
We study three major LLMs compression paradigms, distillation, quantization, and pruning, making our scope comprehensive enough for investigating the effects of diverse compression methods. For distillation, we select four R1 distilled models: DeepSeek-R1-Distill Llama-70B, Qwen-32B, Llama-8B, and Qwen-7B. For quantization, we select 2.51-, 1.73-, and 1.58-bit models by Unsloth [^2] https://huggingface.co/unsloth/DeepSeek-R1-GGUF as the choices of quantized R1 due to their popularity. We also evaluate AWQ Lin et al. (2024), GPTQ (Frantar et al., 2022), GPTAQ (Li et al., 2025), and ANY4/3 (Elhoushi & Johnson, 2025) as reproducible state-of-the-art quantization methods designed for relatively smaller LLMs (e.g., the R1 distilled models). Specifically, we use all four methods to perform 4-bit quantization, and use GPTQ, GPTAQ, and ANY3 for 3-bit quantization as well, since many AWQ implementations do not support 3-bit. For pruning, we run SparseGPT Frantar & Alistarh (2023) on our two largest distilled models. We run interpretaion analysis on linear modules of all layers within LRMs.
2.5 Evaluation Setup
We select four reasoning datasets with varying levels of difficulty: AIME 2024 (Mathematical Association of America, ) for mathematical reasoning, FOLIO (Han et al., 2024) for logical reasoning, Temporal Sequences of BIG-Bench Hard (Suzgun et al., 2022) for temporal reasoning, and MuSiQue (Trivedi et al., 2022) for multihop reasoning. Since MuSiQue requires knowledge memorization besides multihop reasoning, we follow a closed-book setting (directly prompting LRMs to get final answers) to evaluate both reasoning and knowledge retention capabilities. Additional details of benchmarks, along with Table 5 that shows their statistics, are specified in Appendix C.
Accuracy metric is used for AIME 2024, FOLIO, and Temporal Sequences. We adopt exact match (EM) and F1 for MuSiQue. For each model (except R1 and those dynamically quantized LRMs), we run it three times and report its average scores to mitigate performance variability. Implementation details are in Appendix D.
3 Compression Effects on Reasoning Performance
3.1 Overall Performance
The overall performance of R1 and its compressed variants are in Table 1. We show the performance of pruned R1-Distill-Llama-70B and R1-Distill-Qwen-32B under 50% sparsity in Table 1, as it is the default sparsity level of current works (Zhang et al., 2024a; Sun et al., 2023). Additional analysis of test-time compute and model collapse behavior are in Appendices G and H.
Table 1: Benchmark performance of R1 and its compressed variants. All four benchmark scores are averaged over three passes, except the rows marked with . Avg denotes the average scores shown in AIME 2024, FOLIO, and Temporal columns. We segment this table based on model families and mark the highest scores within each model family in bold.